OpenStudy (anonymous):

a woman stands on the bank of a frozen lake with a dog by her side. she skims a bone across the ice at a speed of 3ms. the bone slows down with deceleration 0.4 ms square and the dog chases it with acce;eration 0.6 ms square. how far out from the bank does the dog catch up with the bone

OpenStudy (tylerd):

\[x=3\frac{ m }{ s } \Delta t + \frac{ 1 }{ 2 }(-0.4\frac{ m }{ s^2 }) \Delta t^2\] \[x= \frac{ 1 }{ 2 }(0.6\frac{ m }{ s^2 }) \Delta t^2\]

OpenStudy (tylerd):

\[\frac{ 1 }{ 2 }(0.6\frac{ m }{ s^2 }) \Delta t^2=3\frac{ m }{ s } \Delta t + \frac{ 1 }{ 2 }(-0.4\frac{ m }{ s^2 }) \Delta t^2\]

OpenStudy (tylerd):

if we multiply both sides by 2 we get \[0.6\frac{ m }{ s^2 } \Delta t^2 = 6\frac{ m }{ s } \Delta t -0.4\frac{ m }{ s^2 } \Delta t^2\] then we have to solve for t

OpenStudy (tylerd):

from here we can actually just say \[0.6t^2=6t-0.4t^2\] divide t on both sides we get \[0.6t=6-0.4t\] solve for t \[t=6\]

OpenStudy (tylerd):

plug in 6 for t in both of these equations, to see if they are equal. and x=the distance when the dog catches up. \[3\frac{ m }{ s } \Delta t + \frac{ 1 }{ 2 }(-0.4\frac{ m }{ s^2 }) \Delta t^2\] \[ \frac{ 1 }{ 2 }(0.6\frac{ m }{ s^2 }) \Delta t^2\]